OpenStudy (anonymous):

Trying to setup quadratic equation, but not clear on multiplying these two roots together, I'm not sure if I clear the fractions or just multiply with the fractions? (x-(1/2 + 1/4 Sqrt5))(x-(1/2 -1/4 Sqrt5)) What I am starting with is.. (x-(3/4 Sqrt5)(x-(1/4 Sqrt5) ? any ideas?

OpenStudy (anonymous):

your case is the first one:\[(x-a)(x-b) = x^2 - (a+b)x + ab\]\[(x+a)(x+b) = x^2 + (a+b)x + ab\]\[(x+a)(x-b) = x^2 + (a-b)x - ab\]

OpenStudy (anonymous):

your ab will look like (y + zsqrt(5)) no need to simplify fractions before expanding brackets

OpenStudy (anonymous):

actually there won't be sqrt(5) in ab since they end up canceling out

OpenStudy (anonymous):

I'm not sure I follow?

OpenStudy (anonymous):

So how would they cancel out?

OpenStudy (anonymous):

by adding/multiplying. i did it on paper, scanning it and uploading it ^_^

OpenStudy (anonymous):

ok. great

OpenStudy (anonymous):

i followed my "formula" which is what you'd get by doing FOIL. it avoids a few steps though let me know if you have questions http://img199.imageshack.us/img199/7948/gedd.jpg

OpenStudy (anonymous):

ok. well that seems to work out, I don't know what the answer is supposed to be, but it makes sense. It's a little weird to see it worked out this way. but thank you very much. It'll take me a minute to really understand this. Thank you

OpenStudy (anonymous):

no problem ^_^ to confirm answers, this site is great: http://www.wolframalpha.com/input/?i=%28x-%281%2F2+%2B+1%2F4+Sqrt5%29%29%28x-%281%2F2+-1%2F4+Sqrt5%29%29&dataset=

OpenStudy (anonymous):

you'll wanna scroll down to "alternate forms"

OpenStudy (anonymous):

\[[x-(1+\sqrt{2i})/3][x-(1-\sqrt{2i)/3]}\]

OpenStudy (anonymous):

its pretty unclear what the fraction is: \[\frac{ 1 \pm \sqrt{2i} }{ 3 }\] or\[1 \pm \frac{ \sqrt{2i} }{ 3 }\]

OpenStudy (anonymous):

oh yes. sorry its the first example you wrote

OpenStudy (anonymous):

do I have to multiply by 3 to clear the fraction?

OpenStudy (anonymous):

nope. you can always just add/multiply fractions

OpenStudy (anonymous):

so I would get x^2-1/3+sqrt2i/3-1/3-sqrt2i/3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

let me know if you have questions. sorry for scribble at bottom. ignore the FOIL thing if you already know about it. i can improve my explanation of it but i think you might already know about it: http://img818.imageshack.us/img818/7726/7hvn.jpg

OpenStudy (anonymous):

thank you for your help. I really appreciate it.